OpenStudy (anonymous):

The sides of a square are five to the power of two fifths inches long. What is the area of the square? five to the power of four fifths square inches five to the power of four twenty fifths square inches twenty five to the power of four fifths square inches twenty five to the power of four twenty fifths square inches

OpenStudy (anonymous):

@DanJS @dan815 @ali2x2

OpenStudy (anonymous):

A = s² \[A = 5^\frac{ 2 }{ 5 }*5^\frac{ 2 }{ 5 }\] When you multiply numbers with the same base, add the exponents

OpenStudy (anonymous):

i got A?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay can u help me with one more

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Samantha puts an initial $500 into a savings account. The account has a 4% annual compound interest rate. What is the function that best represents Samantha's account and how much does she have in savings after 5 years? Round your answer to the nearest dollar. Hint: Use the formula, f(x) = P(1 + r)x. f(x) = 500(1.4)x, $2,689 f(x) = 500(0.04)x, $512 f(x) = 500(1.04)x, $608 f(x) = 500(1.04)x, $527

OpenStudy (anonymous):

P is the initial amount, 500. r is the interest rate in decimal, 0.04, so the foumula is \[P=500(1+0.04)^x\] \[P=500(1.04)^x\] Plug in 5 for x to get the amount after 5 years

OpenStudy (anonymous):

@ali2x2 was i right

OpenStudy (ali2x2):

heres a google search im kinda busy http://openstudy.com/study#/updates/55a84f87e4b071e6530c0bfd

OpenStudy (anonymous):

@peachpi i got A am i right